OpenStudy (anonymous):

-3≤2(x+4)<4 sketch the solution set

OpenStudy (anonymous):

how would i start

OpenStudy (anonymous):

your purpose is putting this inequation in order by x so at first, divide all side by 2 and subtract 4

OpenStudy (anonymous):

how do i divide by 2

OpenStudy (anonymous):

-3≤2(x+4)<4 -3/2 ≤ x+4 < 2

OpenStudy (anonymous):

so what do i do next

OpenStudy (anonymous):

and subtract 4 at all side

OpenStudy (anonymous):

how do u subtract 4 from -3/2

OpenStudy (anonymous):

is it -7/2

OpenStudy (anonymous):

Nope, \[-\frac{ 3 }{ 2 } - 4 = -\frac{ 3 }{ 2 } -\frac{ 8 }{ 2 } = -\frac{ 11 }{ 2 }\] do you understand?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

what do u do next

OpenStudy (anonymous):

-3≤2(x+4)<4 -3/2 ≤ x+4 < 2 -11/4 ≤ x < -2 it ends

OpenStudy (anonymous):

thanks